# The reliable quantum master equation of the Unruh-DeWitt detector

### v2

@@ Abstract
In this paper, we present a method for estimating the validity range of the quantum Markovian master equation as applied to the Unruh-DeWitt (UDW) detector within a broader context, particularly without necessitating an exact solution for the detector’s evolution. We propose a relaxed van Hove limit (i.e., late-time limit) and offer a perturbative estimate of the error order resulting from the standard derivation procedure of open quantum dynamics. Our primary findings include reliability criteria for the Markov approximation and conditions for the applicability of the rotating wave approximation (RWA). Nevertheless, the specific forms of these validity conditions rely on the details of the detector-field system, such as the spacetime background, the trajectory of the detector, and the type of quantum field being analyzed. Finally, we illustrate our results by re-examining the open dynamics of an accelerating UDW detector undergoing the Unruh effect, where the validity conditions narrow the parameter space to ensure the solution’s reliability regarding the quantum Markovian master equation.
@@ I Introduction
The particle content of a quantum field is an observer-dependent notion. For example, even in the Poincaré-invariant Minkowski vacuum, an observer with constant acceleration perceives a thermal spectrum sec1-1 . This celebrated ”Unruh effect” can best be manifested by some particle-like quantum probes, among which the best-known idealized model is the Unruh-DeWitt (UDW) detector sec1-2 ; sec1-3 . Formally, it consists of a two-level particle that couples to the background quantum field in a way similar to light-matter interactions while neglecting angular momentum exchange. In a weak coupling limit, i.e., assuming the coupling constant is $g\ll 1$, the response to the detector-field interaction $H_{\text{int}}$ can be solved perturbatively. For an accelerating detector in Minkowski spacetime, its asymptotic transition rate from the ground state to the excited state exhibits a thermal Planckian spectrum with a temperature proportional to its acceleration. This is regarded as a result of detailed-balance between the absorption and emission of field quanta, thus manifesting the particle content of the background quantum field sec1-4-1 ; sec1-4-2 . Much work has been carried out to examine the response and thermal behavior of a detector in various contexts sec1-5 , such as black holes sec1-6 ; sec1-7 , expanding universe sec1-8 ; sec1-9 , quantum gravity sec1-10 , etc.
@@
An alternative perspective is to treat the UDW detector within the framework of the theory of open quantum systems sec1-12 . The time evolution of the detector as a local open system is then governed by a quantum Markovian master equation (QMME), where its interaction with background quantum fluctuations acts as the ”bath” inducing dissipation and decoherence terms. In his seminal works sec1-13 ; sec1-14 , Davies presented a rigorous mathematical proof demonstrating that, once the bath meets general conditions, the QMME for open dynamics becomes accurate in the van Hove limit (i.e., $g^{2}t\sim\mathcal{O}(1)$ while taking $g\rightarrow 0$ and $t\rightarrow\infty$ simultaneously), as all higher-order terms tend toward zero. Consequently, the open system approach is particularly powerful for tracking long-time processes, such as thermalization. In recent years, intensive works sec1-17 ; sec1-18 ; sec1-19 ; sec1-20 ; sec1-21 ; sec1-22 ; sec1-23 ; sec1-24 ; sec1-24+1 have been undertaken by applying this method to the UDW detector in various spacetimes, whose complete dynamics encode the thermal nature of Hawking-Unruh-type effects. For multi-UDW detectors, this method highlights the possibility for generating non-classical correlations, which stem from collective open dynamics and can be observed through various correlation witnesses sec1-25 ; sec1-26 ; sec1-27 ; sec1-28 ; sec1-29 ; sec1-30 , such as entanglement monotones, nonlocality, and quantum uncertainty bounds, depending on detector trajectories, inter-distance between detectors, the types of fields in different spacetimes, etc.
@@
Despite its productive applications, the relevance of the master equation regarding the intricate interactions between UDW detectors and quantum fields is frequently overlooked. In fact, most studies consider the QMME a ”black box” and accept it as valid without further justification. This oversight partly arises from the challenge of assessing the validity of the QMME by directly comparing its solution to the exact one. Specifically, the precise open dynamics are shaped by a differential-integral equation, whose solution at a specific moment relies on its entire evolutionary history and is, therefore, formidable. Conversely, the QMME can establish a quantum dynamical map (e.g., in the form of the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation sec1-15 ; sec1-16 ) that connects open system states at arbitrary times, just by suitably disregarding the aforementioned memory effect through several approximations:
@@
$\bullet$ Born approximation. In cases of weak coupling between the detector and quantum fields, it can be approximated that the state of the field remains unchanged throughout its interaction with the detector. This assumption serves to simplify the structure of the exact evolution equation; however, it remains challenging to solve due to its dependence on the entire history of the detector’s evolution.
@@
$\bullet$ Markov approximation. Assuming that the correlation function of the quantum field $\mathcal{W}(t)$ decays with a characteristic timescale $t_{B}$, e.g., $\mathcal{W}(t)\sim e^{-t/t_{B}}$, once the typical evolution timescale of the detector is much larger than the correlation time (i.e., $t\gg t_{B}$), the effective integration region over time falls within $t\in\left[0,t_{B}\right)$. Consequently, the detector’s evolution at any specific time relies solely on its state at that moment, indicating that the memory effect of the equation has been neglected.
@@
$\bullet$ Rotating wave approximation (RWA). The result of the ”Born-Markov” approximation is a non-CP (completely positive) evolution equation known as the Redfield equation, indicating that the density matrix is notoriously mapped to non-positive matrix. This dilemma underscores the necessity of further utilizing the RWA: all rapidly oscillating terms relative to the typical evolution timescale in the Redfield equation have a negligible effect on detector evolution and can, therefore, be disregarded. While the RWA further limits the domain of the master equation, the combined ”Born-Markov-RWA” approximation consistently ensures that the final derived QMME remains completely positive definite sec1-12 .
@@
In this paper, we reassess the validity of the quantum Markovian master equation within a broader context, specifically without requiring the precise solution of the detector. We invoke the Born approximation to exclude the backreaction of the background quantum field, which is reasonable since it possesses infinite degrees of freedom spreading throughout spacetime. Instead, we address and estimate the primary error introduced by neglecting the memory effect via the Markov approximation. Assuming a small nonvanishing $g$ for the UDW detector model, we propose a relaxed van Hove limit (we refer to late-time limit $g^{2}t\sim\mathcal{O}(1)$). By expanding the detector state in terms of its evolution history, which corresponds to a perturbative series of powers of $g$, we find that the first-order memory effect results in an error of order $g^{4}$. Consequently, the $g$-dependent reliability conditions (i.e., Eqs.(31)) for the Markov approximation can be established by requiring that the first-order memory effect is negligible. Under the late-time limit, we can further identify from the reliability conditions the timescale at which the master equation becomes reliable for the UDW detector-field interaction. On the other hand, unlike previous studies sec1-33 ; sec1-34 ; sec1-35 ; sec1-36 ; sec1-37 , after closely revisiting Davies’ theorems on the rigorous derivation of the QMME, we show that the RWA remains necessary for restoring the CP property of the UDW detector evolution. However, some applicability conditions (i.e., Eqs.(48)) should primarily be satisfied.
The structure of this paper is as follows: In Section II, we review the Unruh-DeWitt detector model and the derivation of its QMME. In Section III, we determine the valid range of the QMME, which is summarized to two sets of conditions. Finally, in Section IV, we re-examine the Unruh effect as an application of our results and determine the range of physical parameters to guarantee the validity of the QMME prediction. For simplicity, throughout the paper, we use natural units $\hbar=c=k_{B}=1$.
@@ II The dynamic equation for the Unruh-DeWitt detector
We formulate the open dynamic equation of the UDW detector and examine how the QMME arises from three approximations sec2-1 : the Born, Markov, and RWA approximations.
@@ II.1 The setup
A UDW detector functions as a two-level quantum system interacting with a quantum field in fixed spacetime. The detector evolves on a classical trajectory due to this interaction. The Hamiltonian of the detector-field system takes the form:
@@
Working within the interaction picture, the momentum and field operators are transformed, like:
@@
with $\mathcal{U}_{\alpha}=\exp\left(-iH_{d}t\right)$ and $\mathcal{U}_{\phi}=\exp\left(-iH_{\phi}t\right)$, respectively. Then, the interaction Hamiltonian becomes:
@@
Let $\rho(t)$ represent the total density matrix of the detector-field system. The evolution of $\rho(t)$ is governed by the Liouville-von Neumann equation:
@@
Since only the time evolution of the UDW detector is relevant, we can trace over all field degrees of freedom from the total state (5) to derive the dynamic equation for the detector density matrix $\rho_{d}(t)=\operatorname{Tr}_{\Phi}(\rho)$ as:
@@
Changing variables to $\tau=t-s$, so that $\int_{0}^{t}ds=-\int_{t}^{0}(-d\tau)=\int_{0}^{t}d\tau$, Eq. (6) can be rewritten as:
@@ II.2 Derivation of quantum Markovian master equation
In a ”bath” (quantum field) with significantly larger degrees than the detector, it is reasonable to assume that the detector evolves while the field remains largely unaffected. Thus, the state of the detector-field system at time $t$ takes on a product form as
@@
Under this Born approximation, Eq. (7) becomes:
@@
The first term on the right-hand side can be reduced to:
@@
We aim to derive a time-local (i.e., memoryless) differential equation for $\rho_{d}$, which depends solely on $t$ and not on the detector’s history. To accomplish this, we assume the background field has a short correlation time $\tau_{B}$, meaning the corresponding Wightman function decays rapidly, such as e.g., $\left|\mathcal{W}(\tau)\right|\sim e^{-\tau/\tau_{B}}$, and becomes zero for $\tau\gg\tau_{B}$. Under this Markov approximation, the main contribution to the integral in Eq.(11) comes from the vicinity of $\tau\approx 0$. The short ”memory” of the bath correlation function allows us to replace $\rho_{d}(t-\tau)\approx\rho_{d}(t)$ without introducing significant errors, which results in the Redfield equation:
$\dot{\rho}_{d}=g^{2}\int_{0}^{\infty}d\tau\left\{\mathcal{W}(\tau)\left[\mathfrak{M}(t-\tau)\rho_{d}(t),\mathfrak{M}(t)\right]+h.c.\right\}.$ (13)
@@
Unfortunately, the Redfield equation is unreliable as it fails to maintain the CP (completely positive) property of quantum dynamics, risking a non-positive density matrix sec2-2 . One may argue that the problem stems from an improperly performed Markovian limit. Recently, it was proposed sec1-33 ; sec1-37 that an alternative Markov approximation, approximating both system state $\rho_{d}(t-\tau)$ and observables $\mathfrak{M}(t-\tau)$ as memoryless. Essentially, their alternative approach can be understood as a synergistic combination of the van Hove limit and the singular coupling limit plus-alicki , where the latter effectively replaces $\mathcal{W}(\tau)$ in Eq.(13) with a delta-function like $\mathcal{W}(\tau)\approx\Lambda\delta(\tau)$, a treatment primarily applied in the high-temperature regime. This leads us to a different master equation, formally as
@@
To attain this, one first converts the monopole momentum operator to the frequency domain:
@@
where $\mathfrak{M}_{\omega}:=\sigma_{+}$ and $\mathfrak{M}_{-\omega}:=\sigma_{-}$. The two terms in the commutator $\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\rho_{d}(t)\right]$ become:
@@
$\dot{\rho}_{d}=g^{2}\sum_{\alpha,\alpha^{\prime}}\left\{\Gamma(\alpha)e^{i\left(\alpha^{\prime}-\alpha\right)t}\left[\mathfrak{M}_{\alpha}\rho_{d}(t),\mathfrak{M}_{\alpha^{\prime}}^{\dagger}\right]+h.c.\right\}.$ (17)
@@
For later convenience, it is important to note sec1-13 ; sec1-14 that the function $\Gamma(\alpha)$ can be separated into the real and imaginary parts
@@
where $\gamma(\alpha)$ is the Fourier transform of bath correlator
@@
In a integral form, the Eq.(17) can be written as:
$\rho_{d}(t)=-g^{2}\int_{0}^{t}d\tau\left(\sum_{\alpha}\Gamma(\alpha)\left[m_{\alpha}^{+},m_{\alpha}\rho_{d}(\tau)\right]+\sum_{\alpha\neq\alpha^{\prime}}\Gamma(\alpha)e^{i\left(\alpha^{\prime}-\alpha\right)t}\left[m_{\alpha^{\prime}}^{+},m_{\alpha}\rho_{\alpha}(\tau)\right]+h.c.\right),$ (22)
where the summation has been split into two parts. One can observe that when $t\gg\left|\alpha^{\prime}-\alpha\right|^{-1}$, the terms with summation over $\alpha\neq\alpha^{\prime}$ oscillate so rapidly that they average to zero. This captures the essence of RWA. Recall that the Markov master equation is valid for $t\gg\tau_{B}$. Likewise, as we examine a longer time scale, the oscillating terms in Eq. (22) will appear increasingly ”fast-oscillating”, which makes the RWA more accurate.
Eventually, by introducing Eq.(19) and the RWA into the Redfield equation, we arrive at the so-called Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) form of QMME as sec1-15 ; sec1-16 :

### v3

# Reliable quantum master equation of the Unruh-DeWitt detector

@@ Abstract
In this paper, we present a method for estimating the validity range of the quantum Markovian master equation as applied to the Unruh-DeWitt (UDW) detector within a broader context, particularly without necessitating an exact solution for the detector’s evolution. We propose a relaxed van Hove limit (i.e., late-time limit) and offer a perturbative estimate of the error order resulting from the standard derivation procedure of open quantum dynamics. Our primary findings include reliability criteria for the Markov approximation and conditions for the applicability of the rotating wave approximation. Nevertheless, the specific forms of these validity conditions rely on the details of the detector-field system, such as the spacetime background, the trajectory of the detector, and the type of quantum field being analyzed. Finally, we illustrate our results by reexamining the open dynamics of an accelerating UDW detector undergoing the Unruh effect, where the validity conditions narrow the parameter space to ensure the solution’s reliability regarding the quantum Markovian master equation.
@@ I Introduction
The particle content of a quantum field is an observer-dependent notion. For example, even in the Poincaré-invariant Minkowski vacuum, an observer with constant acceleration perceives a thermal spectrum sec1-1 . This celebrated ”Unruh effect” can best be manifested by some particlelike quantum probes, among which the best-known idealized model is the Unruh-DeWitt (UDW) detector sec1-2 ; sec1-3 . Formally, it consists of a two-level particle that couples to the background quantum field in a way similar to light-matter interactions while neglecting angular momentum exchange. In a weak coupling limit, i.e., assuming the coupling constant is $g\ll 1$, the response to the detector-field interaction $H_{\text{int}}$ can be solved perturbatively. For an accelerating detector in Minkowski spacetime, its asymptotic transition rate from the ground state to the excited state exhibits a thermal Planckian spectrum with a temperature proportional to its acceleration. This is regarded as a result of detailed balance between the absorption and emission of field quanta, thus manifesting the particle content of the background quantum field sec1-4-1 ; sec1-4-2 . Much work has been carried out to examine the response and thermal behavior of a detector in various contexts sec1-5 , such as black holes sec1-6 ; sec1-7 , expanding universe sec1-8 ; sec1-9 , quantum gravity sec1-10 , etc.
@@
An alternative perspective is to treat the UDW detector within the framework of the theory of open quantum systems sec1-12 . The time evolution of the detector as a local open system is then governed by a quantum Markovian master equation (QMME), where its interaction with background quantum fluctuations acts as the ”bath” inducing dissipation and decoherence terms. In his seminal works sec1-13 ; sec1-14 , Davies presented a rigorous mathematical proof demonstrating that, once the bath meets general conditions, the QMME for open dynamics becomes accurate in the van Hove limit [i.e., $g^{2}t\sim\mathcal{O}(1)$ while taking $g\rightarrow 0$ and $t\rightarrow\infty$ simultaneously], as all higher-order terms tend toward zero. Consequently, the open system approach is particularly powerful for tracking long-time processes, such as thermalization. In recent years, intensive works sec1-17 ; sec1-18 ; sec1-19 ; sec1-20 ; sec1-21 ; sec1-22 ; sec1-23 ; sec1-24 ; sec1-24+1 have been undertaken by applying this method to the UDW detector in various spacetimes, whose complete dynamics encode the thermal nature of Hawking-Unruh-type effects. For multi-UDW detectors, this method highlights the possibility for generating nonclassical correlations, which stem from collective open dynamics and can be observed through various correlation witnesses sec1-25 ; sec1-26 ; sec1-27 ; sec1-28 ; sec1-29 ; sec1-30 , such as entanglement monotones, nonlocality, and quantum uncertainty bounds, depending on detector trajectories, interdistance between detectors, the types of fields in different spacetimes, etc.
@@
Despite its productive applications, the relevance of the master equation regarding the intricate interactions between UDW detectors and quantum fields is frequently overlooked. In fact, most studies consider the QMME a “black box” and accept it as valid without further justification. This oversight partly arises from the challenge of assessing the validity of the QMME by directly comparing its solution to the exact one. Specifically, the precise open dynamics are shaped by a differential-integral equation, whose solution at a specific moment relies on its entire evolutionary history and is, therefore, formidable. Conversely, the QMME can establish a quantum dynamical map (e.g., in the form of the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation sec1-15 ; sec1-16 ) that connects open system states at arbitrary times, just by suitably disregarding the aforementioned memory effect through several approximations:
@@
$(i)$ Born approximation. In cases of weak coupling between the detector and quantum fields, it can be approximated that the state of the field remains unchanged throughout its interaction with the detector. This assumption serves to simplify the structure of the exact evolution equation; however, it remains challenging to solve due to its dependence on the entire history of the detector’s evolution.
@@
$(ii)$ Markov approximation. Assuming that the correlation function of the quantum field $\mathcal{W}(t)$ decays with a characteristic timescale $t_{B}$, e.g., $\mathcal{W}(t)\sim e^{-t/t_{B}}$, once the typical evolution timescale of the detector is much larger than the correlation time (i.e., $t\gg t_{B}$), the effective integration region over time falls within $t\in\left[0,t_{B}\right)$. Consequently, the detector’s evolution at any specific time relies solely on its state at that moment, indicating that the memory effect of the equation has been neglected.
@@
$(iii)$ Rotating wave approximation (RWA). The result of the “Born-Markov” approximation is a non-CP (completely positive) evolution equation known as the Redfield equation, indicating that the density matrix is notoriously mapped to nonpositive matrix. This dilemma underscores the necessity of further utilizing the RWA; all rapidly oscillating terms relative to the typical evolution timescale in the Redfield equation have a negligible effect on detector evolution and can, therefore, be disregarded. While the RWA further limits the domain of the master equation, the combined “Born-Markov-RWA” approximation consistently ensures that the final derived QMME remains completely positive definite sec1-12 .
@@
In this paper, we reassess the validity of the quantum Markovian master equation within a broader context, specifically without requiring the precise solution of the detector. We invoke the Born approximation to exclude the backreaction of the background quantum field, which is reasonable since it possesses infinite degrees of freedom spreading throughout spacetime. Instead, we address and estimate the primary error introduced by neglecting the memory effect via the Markov approximation. Assuming a small nonvanishing $g$ for the UDW detector model, we propose a relaxed van Hove limit [we refer to late-time limit $g^{2}t\sim\mathcal{O}(1)$]. By expanding the detector state in terms of its evolution history, which corresponds to a perturbative series of powers of $g$, we find that the first-order memory effect results in an error of order $g^{4}$. Consequently, the $g$-dependent reliability conditions [i.e., Eqs.(31)] for the Markov approximation can be established by requiring that the first-order memory effect is negligible. Under the late-time limit, we can further identify from the reliability conditions the timescale at which the master equation becomes reliable for the UDW detector-field interaction. On the other hand, unlike previous studies sec1-33 ; sec1-34 ; sec1-35 ; sec1-36 ; sec1-37 , after closely revisiting Davies’ theorems on the rigorous derivation of the QMME, we show that the RWA remains necessary for restoring the CP property of the UDW detector evolution. However, some applicability conditions [i.e., Eqs.(48)] should primarily be satisfied.
The structure of this paper is as follows: In Sec. II, we review the Unruh-DeWitt detector model and the derivation of its QMME. In Sec. III, we determine the valid range of the QMME, which is summarized to two sets of conditions. Finally, in Sec. IV, we reexamine the Unruh effect as an application of our results and determine the range of physical parameters to guarantee the validity of the QMME prediction. For simplicity, throughout the paper, we use natural units $\hbar=c=k_{B}=1$.
@@ II The dynamic equation for the Unruh-DeWitt detector
We formulate the open dynamic equation of the UDW detector and examine how the QMME arises from three approximations sec2-1 ; the Born, Markov, and RWA approximations.
@@ II.1 The setup
A UDW detector functions as a two-level quantum system interacting with a quantum field in fixed spacetime. The detector evolves on a classical trajectory due to this interaction. The Hamiltonian of the detector-field system takes the form,
@@
Working within the interaction picture, the momentum and field operators are transformed, like,
@@
with $\mathcal{U}_{\alpha}=\exp\left(-iH_{d}t\right)$ and $\mathcal{U}_{\phi}=\exp\left(-iH_{\phi}t\right)$, respectively. Then, the interaction Hamiltonian becomes
@@
Let $\rho(t)$ represent the total density matrix of the detector-field system. The evolution of $\rho(t)$ is governed by the Liouville-von Neumann equation,
@@
Since only the time evolution of the UDW detector is relevant, we can trace over all field degrees of freedom from the total state (5) to derive the dynamic equation for the detector density matrix $\rho_{d}(t)=\operatorname{Tr}_{\Phi}(\rho)$ as
@@
Changing variables to $\tau=t-s$, so that $\int_{0}^{t}ds=-\int_{t}^{0}(-d\tau)=\int_{0}^{t}d\tau$, Eq. (6) can be rewritten as
@@ II.2 Derivation of quantum Markovian master equation
In a “bath” (quantum field) with significantly larger degrees than the detector, it is reasonable to assume that the detector evolves while the field remains largely unaffected. Thus, the state of the detector-field system at time $t$ takes on a product form as
@@
Under this Born approximation, Eq. (7) becomes
@@
The first term on the right-hand side can be reduced to
@@
We aim to derive a time-local (i.e., memoryless) differential equation for $\rho_{d}$, which depends solely on $t$ and not on the detector’s history. To accomplish this, we assume the background field has a short correlation time $\tau_{B}$, meaning the corresponding Wightman function decays rapidly, such as e.g., $\left|\mathcal{W}(\tau)\right|\sim e^{-\tau/\tau_{B}}$, and becomes zero for $\tau\gg\tau_{B}$. Under this Markov approximation, the main contribution to the integral in Eq.(11) comes from the vicinity of $\tau\approx 0$. The short “memory” of the bath correlation function allows us to replace $\rho_{d}(t-\tau)\approx\rho_{d}(t)$ without introducing significant errors, which results in the Redfield equation,
$\dot{\rho}_{d}=g^{2}\int_{0}^{\infty}d\tau\left\{\mathcal{W}(\tau)\left[\mathfrak{M}(t-\tau)\rho_{d}(t),\mathfrak{M}(t)\right]+\text{H . c .}\right\}.$ (13)
@@
Unfortunately, the Redfield equation is unreliable as it fails to maintain the CP (completely positive) property of quantum dynamics, risking a nonpositive density matrix sec2-2 . One may argue that the problem stems from an improperly performed Markovian limit. Recently, it was proposed sec1-33 ; sec1-37 that an alternative Markov approximation, approximating both system state $\rho_{d}(t-\tau)$ and observables $\mathfrak{M}(t-\tau)$ as memoryless. Essentially, their alternative approach can be understood as a synergistic combination of the van Hove limit and the singular coupling limit plus-alicki , where the latter effectively replaces $\mathcal{W}(\tau)$ in Eq.(13) with a delta-function like $\mathcal{W}(\tau)\approx\Lambda\delta(\tau)$, a treatment primarily applied in the high-temperature regime. This leads us to a different master equation, formally as
@@
To attain this, one first converts the monopole momentum operator to the frequency domain,
@@
where $\mathfrak{M}_{\omega}:=\sigma_{+}$ and $\mathfrak{M}_{-\omega}:=\sigma_{-}$. The two terms in the commutator $\left[\mathfrak{M}(t),\mathfrak{M}(t-\tau)\rho_{d}(t)\right]$ become
@@
$\dot{\rho}_{d}=g^{2}\sum_{\alpha,\alpha^{\prime}}\left\{\Gamma(\alpha)e^{i\left(\alpha^{\prime}-\alpha\right)t}\left[\mathfrak{M}_{\alpha}\rho_{d}(t),\mathfrak{M}_{\alpha^{\prime}}^{\dagger}\right]+\text{H . c.}\right\}.$ (17)
@@
For later convenience, it is important to note sec1-13 ; sec1-14 that the function $\Gamma(\alpha)$ can be separated into the real and imaginary parts,
@@
where $\gamma(\alpha)$ is the Fourier transform of bath correlator,
@@
In a integral form, the Eq.(17) can be written as
$\rho_{d}(t)=-g^{2}\int_{0}^{t}d\tau\left(\sum_{\alpha}\Gamma(\alpha)\left[m_{\alpha}^{+},m_{\alpha}\rho_{d}(\tau)\right]+\sum_{\alpha\neq\alpha^{\prime}}\Gamma(\alpha)e^{i\left(\alpha^{\prime}-\alpha\right)t}\left[m_{\alpha^{\prime}}^{+},m_{\alpha}\rho_{\alpha}(\tau)\right]+\text{H.c.}\right),$ (22)
where the summation has been split into two parts. One can observe that when $t\gg\left|\alpha^{\prime}-\alpha\right|^{-1}$, the terms with summation over $\alpha\neq\alpha^{\prime}$ oscillate so rapidly that they average to zero. This captures the essence of RWA. Recall that the Markov master equation is valid for $t\gg\tau_{B}$. Likewise, as we examine a longer timescale, the oscillating terms in Eq. (22) will appear increasingly “fast-oscillating”, which makes the RWA more accurate.
Eventually, by introducing Eq.(19) and the RWA into the Redfield equation, we arrive at the so-called Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) form of QMME as sec1-15 ; sec1-16